OpenStudy (trwatkins1):

What is the value of the expression? 19.2 + (–7 + 3.8) A. 16 B. 24 C. 29 D. 30

OpenStudy (anonymous):

19.2 + (–7 + 3.8) 19.2 + (-3.2); = 12.2 - 3.2; = 16

OpenStudy (anonymous):

do parentheses first which give you -3.2. 19.2 + (-3.2) is the same as 19.2-3.2 which is 16